# A physically modelled selection function for compact binary mergers in the LIGO-Virgo O3 run and beyond

### v1

@@ 2.3 Global fits: mass dependence
$\ln\mathcal{L}=-\mu+\sum_{i}\ln\lambda(d_{Li},m_{1i},m_{2i})\,,$ (12)
$\lambda(d_{L},m_{1},m_{2})=P_{\mathrm{det}}(d_{L},m_{1},m_{2})\,p(d_{L})\,p(m_{1},m_{2})\,N_{\text{tot}}\,,$ (13)
$\mu\equiv N_{\text{exp}}=\iiint\lambda(d_{L},m_{1},m_{2})\,\mathrm{d}m_{1}\,\mathrm{d}m_{2}\,\mathrm{d}d_{L}\,.$ (14)
The triple integral in Eq. (14) could in principle be computed numerically, but at high computational cost. Instead, we obtain an accurate approximation from $N_{\mathrm{exp}}\simeq\sum_{i}P_{\mathrm{det}}(d_{Li},m_{1i},m_{2i})$, evaluating over all injections (both found and missed) as these are already correctly distributed as $p(d_{L})p(m_{1},m_{2})$.
@@ 3 Results and discussion
Figure 4(a) shows the values of $d_{\mathrm{mid}}$ obtained from the fit in terms of the detector’s frame masses, computed for every injection in the GWTC-3 catalog. The dependence on masses is smooth, presenting a local maximum around $6200\,\mathrm{Mpc}$ and $m_{1},m_{2}\sim 83\,\mathrm{M_{\odot}}$ in the source frame, which corresponds to chirp masses of $\sim 72\,\mathrm{M_{\odot}}$, consistent with aLIGO sensitivity to compact binary coalescences [59]. For chirp masses below $\sim 100\,\mathrm{M_{\odot}}$, the horizon distance sensitivity increases with bigger masses, since the signal is stronger. However, the merger also shifts towards lower frequencies, eventually leaving the detector frequency band, thus the horizon distance decreases for chirp masses $\gtrsim 100\,\mathrm{M_{\odot}}$. Such behaviour can be seen in the $d_{\mathrm{mid}}$ values from Fig. 4(b), providing reliability to our results. It also shows how for a given chirp mass, $d_{\mathrm{mid}}$ is bigger for equal masses than for unequal ones.
@@
(b) $d_{\mathrm{mid}}$ over redshifted $M_{c}$ and $\eta$.
Figure 4: $d_{\mathrm{mid}}$ values for the GWTC-3 injetions.
Maximum detection probability is plotted in Fig. 5 (left) in terms of redshifted total mass, presenting an interesting behaviour for bigger masses. It is not clear yet the meaning of the local minimum around $M_{z}\sim 400\,\mathrm{M_{\odot}}$, since there could be many effects playing part that could lead to even a nearby signal not being detected at high mass. However, the number of injections at high masses is much smaller, let alone the number of detected ones, making it difficult to obtain precise results in the large mass limit; statistical uncertainties will be discussed in detail in Sec. 5.
Figure 5: Left: Mass dependence of the maximum detection probability. Right: Probability of detection of every GWTC-3 injection over distance rescaled to $d_{\mathrm{mid}}$.
Lastly, Fig. 5 (right) shows the probability of detection of every GWTC-3 injection using the optimized parameters obtained from the global fit. Here, we plot distance rescaled to $d_{\mathrm{mid}}$ (i.e. removing most of the mass dependence), thus $P_{\mathrm{det}}$ follows a expected sigmoid shape going up to almost 1 for small $d_{L}/d_{\mathrm{mid}}$ values and becoming almost zero for very large distances.
@@ 3.1 Diagnostics: cumulative distributions
In order to check our choices of functional forms for both $d_{\mathrm{mid}}$ and $\varepsilon_{\max}$, and also to validate the resulting fit, we compared the cumulative distribution of the actual found injections in the GWTC-3 dataset with the cumulative distribution function (cdf) of the expected found injections using the inferred selection function inferred, given by the following expression,
@@
which represents the probability of the random variable $\theta$ taking on a value less than or equal to $\theta^{*}$. These plots were made for different variables $\theta$: luminosity distance, symmetric mass ratio, total mass and chirp mass in both the source and the detector frame; although only two of them are included here. Figure 6 shows the cumulative plots of found injections over luminosity distance and chirp mass in the detector’s frame. Both analysis validate our $P_{\mathrm{det}}$ fit, since the predicted found injections match almost perfectly the actual detected ones.
(a) Cumulative found injections over $d_{L}$.
(b) Cumulative found injections over $\mathcal{M}_{z}$.
Figure 6: Cumulative found injections from the GWTC-3 data (orange) and the predicted cdf from Eq. (15) (blue).
Lastly, Kolmogorov-Smirnov (K-S) tests were performed for each one of the cdfs, properly checking how good was the fit. Over $d_{L}$, we obtained a K-S statistic of 0.0026 and a $p$-value of 0.65; similar results for $M_{cz}$, K-S statistic = 0.0044 and $p$-value = 0.098. Overall, all the K-S statistics values obtained were $0.004$ or less, i.e. a maximum deviation $\leq 0.4\%$ of the actual from the predicted distribution, and $p$-values of $0.09$ or larger, i.e. no significant discrepancies. Together, the present findings confirm that our model is able to accurately predict the number of found injections, and thus support the choice of the selection function. A limitation of our procedure is that we are unable to constrain or verify the behaviour of $P_{\mathrm{det}}$ outside the physical range of masses and distances covered by the injection set: in extreme cases (for instance binaries with mass ratio $q$ as small as $\mathcal{O}(0.01)$) our model may yield an unphysical estimate. However, its behaviour may readily be checked within the domain where we are estimating BBH merger rates.
Even though these results show that the fit over the mass and distance degrees of freedom is good, we have ignored any dependence on spins. In fact, while carrying out cdfs and K-S tests over spin variables we found significant discrepancies: this motivates including at least the leading dependence of search sensitivity on component spins, which we describe in the following section.
@@ 4 Inclusion of leading spin effects
Up to this point we are not taking into account spin effects, which would only be reasonable if we assume all black hole spins are small. Further work is needed in order to fit aligned spin dependence to the selection function, as well as distance and masses. The spins of individual black holes within a binary system introduce complexities in the emitted gravitational wave, influencing the amplitude, waveform modulation, and precession of the orbital plane. These spin-induced effects not only contribute to the quality of the observed signals but also impact the accuracy of parameter estimation, effective detection ranges of gravitational wave detectors, and the overall event rates [60, 57, 61]. BBH spins, particularly the effective spin $\chi_{\mathrm{eff}}$, are essential for extracting precise astrophysical information, providing valuable insights into the population properties and formation channels of binary black hole systems [62, 63].
Signals from binaries with $\chi_{\mathrm{eff}}>0$ (i.e. aligned spins) have a higher probability of detection, since their waveforms have a larger number of cycles and a higher duration in the detector sensitive band [26]; thus $\chi_{\mathrm{eff}}$ should have a significant impact on the selection function [27, 64]. Looking at the results from the fit described previously, the distribution of found injections over $\chi_{\mathrm{eff}}$ was investigated for various subsets of the injection sets, verifying that the detection probability is an increasing function of $\chi_{\mathrm{eff}}$. After several considerations, the spin dependence was integrated into $d_{\mathrm{mid}}$ as follows:
@@
$f_{AS}=\chi_{\mathrm{eff}}(c_{1}+c_{11}M_{z})\,,$ (17)
where $\chi_{\mathrm{eff}}$ is the effective inspiral spin parameter, given by
where $\vec{s}_{1,2}$ are the (dimensionless) spins of each black hole and $\hat{L}$ is the direction of the orbital angular momentum [26]. We thus require to fit the coefficients $c_{1}$, $c_{11}$ simultaneously with other parameters.
(b) Maximum detection efficiency $\varepsilon_{\max}$ as a function of redshifted total mass $M_{z}$.
Figure 7: Search sensitivity dependences for the fit including aligned spin effects.
@@
Figure 8: $d_{\mathrm{mid}}$ values for the GWTC-3 injections with spin dependence.
Note that some changes to the likelihood of Eq. (12) are needed to account for the spin dependence $\chi_{\mathrm{eff}}(m_{1},m_{2},s_{1z},s_{2z})$: we add a term accounting for the spins’ PDFs into Eq. (13), and also integrate $N_{\mathrm{exp}}$ over $s_{1z}$ and $s_{2z}$ in Eq. (14).
The inclusion of spin dependence improved the fit considerably, giving much higher values of the likelihood. Figure 7 shows the $d_{\mathrm{mid}}$ values from the new fit, where we can see that the sensitive range increases by up to just above a factor of 2 for high-mass systems with high aligned spin. In Fig. 8(b) we can see the effect of $\chi_{\mathrm{eff}}$ in $d_{\mathrm{mid}}$ as a function of chirp mass: basically systems are detectable at higher distances the more aligned the spins are (the closer $\chi_{\mathrm{eff}}$ is to 1) as expected. By comparison with Fig. 4(b), it seems that at a given $\mathcal{M}_{z}$, variations in $\chi_{\mathrm{eff}}$ have a larger effect on detectability than in $\eta$.
The size of the spin correction introduced by $\exp(f_{AS})$ is displayed in Fig. 7(a). It’s mostly close to 1 for middle values of $\chi_{\mathrm{eff}}$ and small masses, but in the high mass limit and extreme values of the effective spin the correction can reduce $d_{\mathrm{mid}}$ down a factor of $\sim 0.5$ for the most negative values of $\chi_{\mathrm{eff}}$ and increase it a factor of $\sim 2.5$ for $\chi_{\mathrm{eff}}\sim 1$. The spin correction also affected the fit of the maximum detection probability, as evident from Fig. 7(b). It’s likely that the lower values of $\varepsilon_{\max}$ from the previous fit were a way of fitting a mixture of different $d_{\mathrm{mid}}$ values due to different spins at the same component mass point.
Lastly, we perform K-S tests to validate this new model with spin dependence. Figure 9 shows the cumulative found injections in the GWTC-3 catalog over the total redshifted mass and $\chi_{\mathrm{eff}}$, along with the cdfs predicted by the selection function: we obtain K-S statistics and $p$-values of $(0.0043,0.11)$ for $M_{z}$, and $(0.0039,0.19)$ for $\chi_{\mathrm{eff}}$. In general, the K-S statistic values obtained were $0.004$ or less again, and $p$-values of $0.12$ or larger, implying that the spin dependence is well fitted.
(a) Cumulative found injections over $M_{z}$.
(b) Cumulative found injections over $\chi_{\mathrm{eff}}$.
Figure 9: Cumulative found injections from the GWTC-3 data (orange) and the predicted cdf from Eq. (15) (blue) with spin correction.
@@ 5 Statistical uncertainties in $P_{\mathrm{det}}$
In order to quantify the statistical uncertainties in our estimation of the selection function, due to the finite number both of njectioins performed and, particularly, of found injections, we use a bootstrap resampling method. We randomly resampling the observed data $n$ times, allowing for replacements, and fit $P_{\mathrm{det}}$ to each bootstrap set of injections, thus obtaining $n$ sets of optimized fit parameters. In addition to estimating the uncertainty on each fit parameter, bootstrap resampling fit allows us to assess how far uncertainties in the parameters are correlated with each other. Figure 10 illustrates the resulting uncertainty estimates in $d_{\mathrm{mid}}$ and $\varepsilon_{\max}$ over the detected total mass and the effective spin, by plotting $n=100$ bootstrap iterations.
@@
The estimated statistical uncertainty is almost negligible for low masses, however in the high mass limit it becomes significant. Moreover, for a fixed total mass the uncertainty decreases with increasing $\eta$, due to the higher density of found injections systems close to equal mass. These trends are expected, as the number of detected injections in the high (redshifted) mass limit is very low: it will be important to take these uncertainties into account when using the fit and drawing conclusions for population analysis over this mass range.

### v2

@@ 2.3 Global fits: mass dependence
$\ln\mathcal{L}=-\mu+\sum_{i}\ln\lambda(d_{L}^{i},m_{1}^{i},m_{2}^{i})\,,$ (12)
$\lambda(d_{L},m_{1},m_{2})=P_{\mathrm{det}}(d_{L},m_{1},m_{2})\,p(d_{L})\,p(m_{1},m_{2})\,N_{\mathrm{tot}}\,,$ (13)
$\mu\equiv N_{\exp}=\int\!\!\int\!\!\int\!\lambda(d_{L},m_{1},m_{2})\,\mathrm{d}m_{1}\,\mathrm{d}m_{2}\,\mathrm{d}d_{L}\,.$ (14)
The triple integral in (14) could in principle be computed numerically, but at high computational cost. Instead, we obtain an accurate approximation from $N_{\mathrm{exp}}\simeq\sum_{i}P_{\mathrm{det}}(d_{Li},m_{1i},m_{2i})$, evaluating over all injections (both found and missed) as these are already correctly distributed as $p(d_{L})p(m_{1},m_{2})$.
@@ 3 Results and discussion
Figure 4(a) shows the values of $d_{\mathrm{mid}}$ obtained from the fit in terms of the redshifted masses, computed for every injection in the GWTC-3 catalog. The dependence on masses presents a local maximum around $3800\,\mathrm{Mpc}$ for $m_{1z},m_{2z}\gtrsim 150\,\mathrm{M_{\odot}}$, which corresponds to redshifted chirp mass $\sim\!150\,\mathrm{M_{\odot}}$, consistent with aLIGO expected sensitivity to compact binary coalescences [75]. For chirp masses below $\sim\!150\,\mathrm{M_{\odot}}$ the sensitive distance increases with mass, mainly driven by a higher signal amplitude. However, the merger also shifts towards lower frequencies, eventually leaving the detectors’ frequency band, thus the sensitive distance decreases for chirp masses $\gtrsim 150\,\mathrm{M_{\odot}}$. Moreover, higher mass signals have fewer cycles visible in band, thus being more readily mimicked by glitches: an effectively higher background for high-mass templates is therefore expected to contribute to decreasing sensitivity towards high masses. Figure. 4(b) shows behaviour consistent with these expectations, and also indicates a trend for higher (lower) sensitivity to equal (unequal)-mass binaries at a given chirp mass.
@@
(b) $d_{\mathrm{mid}}$ over redshifted $\mathcal{M}$ and $\eta$.
Figure 4: Sensitive distance $d_{\mathrm{mid}}$ values for the O3 BBH injections obtained from the global fit of $P_{\mathrm{det}}$, plotted over redshifted masses $m_{1z},m_{2z}$ (a) and over redshifted chirp mass $\mathcal{M}_{z}$ and symmetric mass ratio $\eta$ (b).
Maximum detection probability is plotted in Fig. 5(a) in terms of redshifted total mass, presenting an interesting behaviour for higher masses. The interpretation of the local minimum around $M_{z}\sim 400\,\mathrm{M_{\odot}}$ is not yet clear, since various different effects (precessing spins, non-dominant signal modes, higher glitch backgrounds) could lead to high-mass signals not being detected, even at very low distances: we further discuss spin effects in Sec. 4. However the number of injections performed at high masses is much smaller, let alone the number detected, making it difficult to obtain precise results in the large mass limit; statistical uncertainties will be discussed in detail in Sec. 5.
(a) $\varepsilon_{\max}$ over total redshifted mass.
(b) $P_{\mathrm{det}}$ over distance rescaled to $d_{\mathrm{mid}}$ and over $\varepsilon_{\max}$.
Figure 5: Mass dependence of the maximum detection probability $\varepsilon_{\max}$ (a) and probability of detection over distance rescaled to $d_{\mathrm{mid}}$, colored by $\varepsilon_{\max}$ (b) for O3 BBH injections. The spread of points in (b) indicating a departure from a single universal function is caused by the mass dependence of $\varepsilon_{\max}$.
Lastly, Fig. 5(b) shows the probability of detection of every O3 BBH injection using the optimized parameters obtained from the global fit. Here, we plot distance rescaled to $d_{\mathrm{mid}}$ (i.e. removing most of the mass dependence), thus $P_{\mathrm{det}}$ follows a expected sigmoid shape going up to almost 1 for small $d_{L}/d_{\mathrm{mid}}$ values and becoming almost zero for very large distances.
@@ 3.1 Diagnostics: cumulative distributions
In order to check our choices of functional forms for both $d_{\mathrm{mid}}$ and $\varepsilon_{\max}$, and also to validate the resulting fit, we compared the cumulative counts and (normalized) distributions (cdfs) of actual found injections in the GWTC-3 dataset over various parameters of interest with the cumulative counts or cdfs that are expected from the inferred selection function. The expected cdf of found injections is given by
@@
i.e. the probability of the parameter $\theta$ for a found injection taking on a value less than or equal to $\theta^{*}$. We carried out comparisons for different variables $\theta$: luminosity distance, symmetric mass ratio, total mass and chirp mass in both the source and the detector frame; for the sake of space, we only illustrate two here. Figure 6 compares the cumulative counts $\sum P_{\mathrm{det}}(\theta_{i})$ of found injections over luminosity distance and redshifted chirp mass, validating our $P_{\mathrm{det}}$ fit since the predicted found injections match almost perfectly the actual detected ones.
(a) Cumulative found injection count over $d_{L}$.
(b) Cumulative found injection count over $\mathcal{M}_{z}$.
Figure 6: Cumulative found injections from the GWTC-3 data (orange) and the predicted cdf (’model’) from Eq. (15) (blue), evaluating the model count at the parameter value of each injection.
Lastly, Kolmogorov-Smirnov (K-S) tests were performed for each of the cdfs to quantify the accuracy of the fit. Over $d_{L}$, we obtained a K-S statistic of 0.0026 and a $p$-value of 0.65; similar results for $\mathcal{M}_{z}$, K-S statistic = 0.0044 and $p$-value = 0.098. Overall, all the K-S statistics values obtained were $0.004$ or less, i.e. a maximum deviation $\leq 0.4\%$ of the actual from the predicted distribution, and $p$-values of $0.09$ or larger, i.e. no significant discrepancies. Together, the present findings confirm that our model is able to accurately predict the number of found injections, and thus support the choice of the selection function. We are, though, unable to constrain or verify the behaviour of $P_{\mathrm{det}}$ outside the physical range of masses and distances covered by the injection set: in extreme cases (for instance binaries with mass ratio $q$ as small as $\mathcal{O}(0.01)$) our model may yield an unphysical estimate. However its behaviour may readily be checked within the domain where, for instance, LVK give BBH merger rates estimates [14].
Even though these results show that the fit over the mass and distance degrees of freedom is good, we have ignored any dependence on spins. In fact, comparing cdfs and performing K-S tests over spin variables shows significant discrepancies: this motivates including at least the leading dependence of search sensitivity on component spins, which we describe in the following section.
@@ 4 Inclusion of leading spin effects
Up to this point we are not taking into account spin effects, which would only be reasonable if we assume all black hole spins are small. Further work is needed in order to fit aligned spin dependence to the selection function, as well as distance and masses. The spins of individual black holes within a binary system introduce complexities in the emitted gravitational wave, influencing the amplitude, waveform modulation, and precession of the orbital plane. These spin-induced effects not only contribute to the quality of the observed signals but also impact the accuracy of parameter estimation, effective detection ranges of gravitational wave detectors, and the overall event rates [76, 72, 77]. BBH spins, particularly the effective spin $\chi_{\mathrm{eff}}$, are essential for extracting precise astrophysical information, providing valuable insights into the population properties and formation channels of binary black hole systems [78, 79]. The effective inspiral spin parameter $\chi_{\mathrm{eff}}$ is defined as [27]
where $\vec{s}_{1,2}$ are the (dimensionless) spins of each black hole and $\hat{L}$ is the direction of the orbital angular momentum; this defines the spin projections along the orbital axis $s_{1z},s_{2z}$.
(b) $\varepsilon_{\max}$ as a function of $M_{z}$.
Figure 7: Search sensitivity dependencies for the fit including aligned spin effects. Panel (a) shows the correction to the sensitive distance, $f_{AS}$, over total redshifted mass and effective spin, for every O3 BBH injection. The comparison between the maximum search efficiency dependence on redshifted total mass from the fit with and without spin correction in shown in (b).
Signals from binaries with $\chi_{\mathrm{eff}}>0$ (i.e. aligned spins) have a higher probability of detection, since their waveforms have a larger number of cycles and a higher duration in the detector sensitive band [27]; thus $\chi_{\mathrm{eff}}$ should have a significant impact on the selection function [28, 80]. As mentioned previously, we compared cdfs of found injections, as well as histogramming them over various parameters, in order to study the dependence of $P_{\mathrm{det}}$ on $\chi_{\mathrm{eff}}$. Significant discrepancies were found with the previous fit function neglecting spins, Eqs. (9-11), verifying that the detection probability is an increasing function of $\chi_{\mathrm{eff}}$, an effect appearing stronger at higher total mass. Taking into account these discrepancies, spin dependence was implemented as follows: [^5] Here the $c$ subscripts represent powers of $M_{z}$ and $\chi_{\mathrm{eff}}$.
@@
$f_{AS}=\chi_{\mathrm{eff}}(c_{01}+c_{11}M_{z})\,.$ (18)
We thus require to fit the coefficients $c_{01}$, $c_{11}$ simultaneously with other parameters. Some changes to the likelihood are needed to account for the spin dependence $\chi_{\mathrm{eff}}(m_{1},m_{2},s_{1z},s_{2z})$: in place of Eqs. (12-14) we have:
$\ln\mathcal{L}=-\mu+\sum_{i}\ln\lambda(d_{L}^{i},m_{1}^{i},m_{2}^{i},s_{1z}^{i},s_{2z}^{i})\,,$ (19)
$\quad\lambda(d_{L},m_{1},m_{2},s_{1z},s_{2z})=P_{\mathrm{det}}(d_{L},m_{1},m_{2},\chi_{\mathrm{eff}})p(d_{L})p(m_{1},m_{2})p(s_{1z},s_{2z})N_{\mathrm{tot}}\,,$ (20)
$\mu\equiv N_{\exp}=\int\!\lambda(d_{L},m_{1},m_{2},s_{1z},s_{2z})\,\mathrm{d}m_{1}\,\mathrm{d}m_{2}\,\mathrm{d}s_{1z}\,\mathrm{d}s_{2z}\,\mathrm{d}d_{L}\,.$ (21)
@@
Figure 8: Sensitive distance $d_{\mathrm{mid}}$ values for the O3 BBH injections from the fit with spin dependence, over redshifted masses (a) and over redshifted chirp mass and effective spin (b).
We find that the spin-dependent factor $\exp(f_{AS})$, shown in Fig. 7(a), is close to 1 for small values of $\chi_{\mathrm{eff}}$ and small masses, but in the high mass limit, extreme values of the effective spin the correction can reduce $d_{\mathrm{mid}}$ by a factor of $\sim 2$ for the most negative values of $\chi_{\mathrm{eff}}$ and increase it by a factor of $\sim 2.5$ for $\chi_{\mathrm{eff}}\sim 1$. The spin correction also affected the fit of the maximum detection probability, as evident from Fig. 7(b). It is plausible that the lower values of $\varepsilon_{\max}$ from the fit neglecting spins were driven by the need to fit a mixture of different $d_{\mathrm{mid}}$ values due to different spins at the same component mass point. The remaining apparent trend in $\varepsilon_{\max}$ is still subject to high uncertainties, explored in Sec. 5.
Figure 8(a) shows the $d_{\mathrm{mid}}$ values in terms of the redshifted masses from the new fit, where we can see that the maximum sensitive range increases by almost a factor of 2. In Fig. 8(b) we can see the effect of $\chi_{\mathrm{eff}}$ in $d_{\mathrm{mid}}$ as a function of chirp mass: basically systems are detectable at higher distances the more aligned the spins are (the closer $\chi_{\mathrm{eff}}$ is to 1) as expected. By comparison with Fig. 4(b), at fixed $\mathcal{M}_{z}$ variations in $\chi_{\mathrm{eff}}$ appear to have a larger effect on detectability than variations in $\eta$.
Lastly, we perform K-S tests to validate this new model with spin dependence. Figure 9 shows the cumulative found injections in the GWTC-3 catalog over $\chi_{\mathrm{eff}}$, along with the cdfs predicted by the selection function, for our fits both without the spin factor (a), and including it (b). A significant mismatch is visible in the fit without spin dependence. In general, the K-S statistic values over different variables obtained with the spin correction were $0.004$ or less again, and $p$-values of $0.12$ or larger, implying that the spin dependence is well fitted.
(a) Cumulative found injections over $\chi_{\mathrm{eff}}$ without
the spin correction.
(b) Cumulative found injections over $\chi_{\mathrm{eff}}$ with
the spin correction.
Figure 9: Comparative of the cumulative found injections over $\chi_{\mathrm{eff}}$ from the GWTC-3 data (orange) and the predicted cdf from (15) (blue), without (a) and with (b) spin correction. The K-S statistics and $p$-values for the model without spin correction (a) are $(0.0438,0)$, and $(0.0039,0.19)$ for the new model (b).
@@ 5 Statistical uncertainties in $P_{\mathrm{det}}$
In order to quantify the statistical uncertainties in our estimation of the selection function, due to the finite number both of injections performed and, particularly, of found injections, we use a bootstrap resampling method. We randomly resample the observed data $n$ times, allowing for replacements, and fit $P_{\mathrm{det}}$ to each bootstrap set of injections, thus obtaining $n$ sets of optimized fit parameters. In addition to estimating the uncertainty on each fit parameter, bootstrap resampling fit allows us to assess how far uncertainties in the parameters are correlated with each other. Figure 10 illustrates the resulting uncertainty estimates in $d_{\mathrm{mid}}$ and $\varepsilon_{\max}$ over the detected total mass and the effective spin, by plotting $n=100$ bootstrap iterations.
@@
The results show that the estimated statistical uncertainty is lowest for low masses, becoming significant in the high mass limit; for a few bootstrap iterations we even see instability in the behaviour at very high total mass (over $\sim\!200\,\mathrm{M_{\odot}}$) for unequal mass binaries, which we attribute to a very small number of found injections in this region. Figure 10(a) additionally shows that at fixed total mass the uncertainty decreases with increasing $\eta$, due to an increasing density of found injections. Figure 10(b), showing the uncertainty in $d_{\mathrm{mid}}$ as a function of effective spin for different values of the redshifted total mass, again demonstrates that uncertainties grow with $M_{z}$, but remain near-constant over $\chi_{\mathrm{eff}}$. Figure 10(c) shows the uncertainty in the dependence of $\varepsilon_{\max}$ on $M_{z}$: while there is a significant decrease from masses $100\,\mathrm{M_{\odot}}$ up to $350-400\,\mathrm{M_{\odot}}$, the behaviour at very high $M_{z}$ is subject to large uncertainties, with a slight (not highly significant) trend to increasing $\varepsilon_{\max}$.
Most of these trends are expected, as the number of detected injections in the high (redshifted) mass limit is very low. Obtaining useful sensitivity estimates in this regime will require a much higher density of injections here, allowing us to diagnose more clearly what factors may be causing the apparent increase in $\varepsilon_{\max}$. It will be important to account for such uncertainties in any population analysis over this mass range; note that the main LVK O4 population analyses did not explore such high masses [14] given the lack of clear detections.